A health clinic uses a solution of bleach to sterilize petri dishes in which cultures are grown. The sterilization tank contains 100 gal of a solution of 2% ordinary household bleach mixed with pure distilled water. New research indicates that the concentration of bleach should be 5% for complete sterilization. How much of the solution should be drained and replaced with bleach to increase the bleach content to the recommended level?
Thank you.

Answer by Paul(988)

0.02(100)+x=0.05(100+x)
2 + x = 5 + 0.05x
Simplify

-0.95x = -3
x = 3.16

Hence, about 3.16 of solution
Paul.